# EgoExo-Fitness: Towards Egocentric and Exocentric Full-Body Action Understanding

### v1

@@ Abstract
We present EgoExo-Fitness, a new full-body action understanding dataset, featuring fitness sequence videos recorded from synchronized egocentric and fixed exocentric (third-person) cameras. Compared with existing full-body action understanding datasets, EgoExo-Fitness not only contains videos from first-person perspectives, but also provides rich annotations. Specifically, two-level temporal boundaries are provided to localize single action videos along with sub-steps of each action. More importantly, EgoExo-Fitness introduces innovative annotations for interpretable action judgement–including technical keypoint verification, natural language comments on action execution, and action quality scores. Combining all of these, EgoExo-Fitness provides new resources to study egocentric and exocentric full-body action understanding across dimensions of “what”, “when”, and “how well”. To facilitate research on egocentric and exocentric full-body action understanding, we construct benchmarks on a suite of tasks (i.e., action classification, action localization, cross-view sequence verification, cross-view skill determination, and a newly proposed task of guidance-based execution verification), together with detailed analysis. Code and data will be available at https://github.com/iSEE-Laboratory/EgoExo-Fitness/tree/main.
@@
Egocentric video dataset Full-body action understanding Fitness practising
@@ 1 Introduction
By looking at the field of egocentric video understanding, we find that egocentric full-body action understanding is yet to be well explored due to the lack of datasets. Existing egocentric video datasets primarily focus on interactive actions like desktop works [36, 10, 11, 12, 57, 21, 2, 68] (e.g., cooking and assembling) and daily interaction [22, 13, 52, 46, 60, 69] (i.e., interacting with daily objects or humans). The other branch of egocentric datasets [33, 77, 72, 24] mainly focuses on body pose estimation and reconstruction rather than understanding full-body action from other dimensions (e.g., verifying the consistency of action sequences and assessing the execution of action).
@@
  • Secondly, it provides two-level temporal boundaries to localize a single fitness action as well as sub-steps (i.e., getting ready, executing, and relaxing) of each action.
  • Lastly, EgoExo-Fitness introduces annotations on interpretable action judgement, including technical keypoint verification, natural language comment on action execution, and quality score for each single action video. To our knowledge, no previous dataset contains such annotations on action judgement.
Combining all of these, EgoExo-Fitness spans 31 hours with 1248 cross-view action sequence videos featuring more than 6000 single fitness actions. With synchronized ego-exo videos and rich annotations, EgoExo-Fitness provides new resources to study egocentric and exocentric full-body action understanding across dimensions of “what”, “when”, and “how well”.
@@
In summary, the contributions of our work are as follows: 1) We present EgoExo-Fitness, a new full-body action understanding dataset featuring fitness sequence videos recorded from synchronized egocentric and exocentric cameras; 2) We introduce rich annotations on EgoExo-Fitness, including two-level temporal boundaries and novel annotations of interpretable action judgement; 3) We construct benchmarks on five relevant vision tasks, including the newly introduced Guidance-based Execution Verification (GEV) and extensive experimental analysis.
@@ 2.1 Revisiting Current Datasets
Full-body action understanding datasets. Human body movements contain complex motion patterns and technical skills, presenting a series of challenges for Full-Body Action Understanding (FBAU). To address these challenges, datasets like NTU-RGB+D[58, 44], Human3.6M [29], Diving48 [40] and FineGym [59] are proposed to enable research on recognizing coarse-and-fine human full-body actions. Beyond recognition, datasets like Diving48-SV [53] and RepCount [26] are present to address tasks (e.g., Sequence Verification and Repetitive Action Counting) that require stronger temporal modeling ability. Note that technical full-body action videos (e.g., diving and vaulting) will reflect human skills. Hence, in recent years, datasets for Action Assessment, like AQA-7[49], FineDiving [71], LOGO[76], are introduced to study the subtle skill differences between action videos. Another branch of datasets [62, 18, 6] focuses on estimating or reconstructing 3D human poses from full-body action videos, achieving the development of Virtual Reality. Though great progress has been achieved, today’s full-body action understanding datasets mainly assume that human full-body action videos are captured by exocentric cameras. Such an assumption limits further exploration in more flexible settings. Moreover, some datasets (e.g., WEAR [5] and 1st-basketball [3]) propose to understand sports and fitness activities from egocentric viewpoints. However, these datasets are limited by their scale and task-specific annotations.
@@
Egocentric video datasets. Egocentric Video Understanding (EVU) has great application value for AR/VR and Robotics. Most existing EVU datasets focus on interactive actions: 1) tabletop activities in kitchen [10, 11, 12, 39, 64] or on a static working platform [36, 57, 21, 2, 55]; 2) daily activities interacting with daily objects [69, 37, 60, 46, 31, 27] or individuals [47, 56, 15]. Although recently proposed Ego4D [22] expands beyond interactive activities to a wider variety of daily activities, works on this branch of datasets rarely focus on egocentric full-body action understanding. Another branch of work aims to estimate or reconstruct full-body pose from egocentric videos, and several datasets [38, 77, 1, 67, 33] are released.
@@
Different from existing datasets, EgoExo-Fitness features synchronized egocentric and exocentric videos of full-body fitness actions and provides rich annotations (especially novel annotations of interpretable action judgement) for future research on understanding ego- and exo-centric full-body actions across dimensions of “what”, “when”, and “how well”.
@@
It is worth noting that a concurrent large dataset, Ego-Exo4D [23], also contains ego-exo full-body (physical) action videos and annotations on how well an action is performed. EgoExo-Fitness still has its values: (1) it focuses on a novel scenario (i.e., natural fitness practicing); (2) it provides novel annotations(e.g., technical keypoints verification), supporting the novel task on interpretable action assessment. We will provide detailed comparisons across our work and Ego-Exo4D in Sec. 3.5 and Sec. A4.2.
@@ 2.2 Revisiting Relevant Tasks
Action Classification & Localization. As the fundamental tasks in video action understanding, action classification [4, 9, 35] and temporal action localization [45, 75, 66] are widely explored in previous work. In our work, we benchmark EgoExo-Fitness on action classification and localization to present the domain gap of views and show the challenges of ego- and exo-centric full-body action understanding.
Sequence Verification. Sequence Verification (SV) [53, 14, 25] is proposed to study the action order of sequential videos under a scenario that precise temporal annotations are not provided. Today’s benchmark on SV rather focuses on exocentric-SV (i.e., COIN-SV and Diving48-SV) or egocentric-SV (i.e., CSV). In this work, we present the first benchmark on cross-view sequence verification and provides extensive experimental analysis.
Action Assessment. Existing datasets on Action Assessment (or Skill Determination) are mainly based on videos from either ego-cameras [16, 20] or exo-cameras [71, 50, 49, 76, 70, 17, 74, 19], which leads to single-view assessment ability. Also, today’s popular AQA datasets only provide the annotations on action scores or pair-rankings, which is unable existing work to directly explore the interpretability of the predicted results. To address the first issue, we introduce the first benchmark on Cross-View Skill Determination. For the second issue, we propose a novel task, Guidance-based Execution Verification (GEV), which aims to verify whether the execution of an action satisfies the given technical keypoints.
@@ 3 EgoExo-Fitness Dataset
In this section, we introduce the EgoExo-Fitness dataset in detail. We will start by describing the recording system in Sec. 3.1, data collection in Sec. 3.2, and annotations in Sec. 3.3. Finally, we present the statistics and comparison with related datasets in Sec. 3.4 and Sec. 3.5, respectively.
@@ 3.1 Recording System
We build a recording system for EgoExo-Fitness to capture action videos from egocentric and exocentric views. Fig. 2 shows the setups of our recording system. For egocentric video capturing, we design a headset with multiple cameras to record videos from forward and downwards views. Specifically, we use a GoPro to record the forward (i.e., Ego-M) view of participants and apply two Insta Go3 to record the left-downward (i.e. Ego-L) and right-downward (i.e., Ego-R) views. For the exocentric cameras, we locate them at the participants’ front (i.e., Exo-M), left-front (i.e., Exo-L), and right-front (i.e., Exo-R) sides and ensure they can record full-body actions completely. All cameras are synchronized manually using a timed event that is visible from them.
@@ 3.2 Action Sequence & Recording Protocols
Following FLAG3D [62] and HuMMan [6], we select 12 types of fitness actions based on various driving muscles (i.e., chest, abdomen, waist, hip, and whole body). All selected actions are listed in Tab. 1. In the remaining part of the paper, for convenience, we will use the abbreviation to present the actions. Furthermore, to enrich the temporal diversity of the recorded videos, we define 76 action sequences by randomly combining 3 to 6 different actions. For example, “starting with Push-ups, then Sit-ups, finally High Knee” is an action sequence with three fitness actions. For details of the action sequences, please refer to Appendix A2.
@@
Table 1: Recorded fieness actions in EgoExo-Fitness. Abbr.: the abbreviation of the fitness action.
Action types Abbr. Action types Abbr.
1: Kneeling Push-ups KPU 7: Leg Reverse Lunge LRL
2: Push-ups PU 8: Leg Lunge with Knee Lift LLKL
3: Kneeling Torso Twist KTT 9: Sumo Squat SS
4: Knee Raise and Abdomi- KRAMC 10: Jumping Jacks JJ
         nal Muscles Contract 11: High Knee HK
5: Shoulder Bridge SB 12: Clap Jacks CJ
6: Sit-ups SU
@@ 3.3 Annotations
Interpretable Action Judgement. Our motivations for providing this series of annotations are two-folds. First, it is easy for human experts to compare an action video and the text guidance to conclude whether the actor followed the guidance or not and point out which technical keypoint in the text guidance is missed during the execution. However, such ability has rarely been studied in existing video-language answering and video-language retrieval works. Second, although Action Assessment [49, 70, 50, 71, 16, 17, 41] has been studied for many years, and great progress has been achieved, interpretable action assessment based on language annotations has never been explored due to the lack of well-collected dataset. Also, existing action assessment work is limited to ego-only or exo-only scenarios due to the collecting manner of the datasets.
To address these issues, we develop a web-based annotation tool for EgoExo-Fitness, and collect three categories of interpretable action judegment annotations (i.e., Technical keypoints Verification, Natural Language Comment, and Action Quality Score) step by step. We will introduce the details as follows.
@@
(1) Technical Keypoints Verification. A paragraph of text guidance on fitness action can be divided into several technical keypoints. By following the keypoints, one can achieve the goal of exercise while avoiding physical injury. In EgoExo-Fitness, we provide the verification annotations on the keypoints for each single action in the following three steps. First, following FLAG3D [62], we provide a paragraph of text guidance for each recorded action. Second, we prompt LLM (e.g., ChatGPT) to separate the text guidance into several keypoints. Third, we ask the annotators to verify an action by comparing the execution with the technical keypoints. Given a single action video and a technical keypoint, if the action satisfies the keypoint, an annotation of “True” will be provided (otherwise “False”).
@@
(3) Action Quality Score. Finally, the annotators are asked to score the actions from 1 to 5 (worst to best) based on the technical keypoint verifications and comments they make.
@@ 3.4 Statistics
Number of recordings and Durations. EgoExo-Fitness collects 1248 cross-view action sequence videos from 76 action sequences, which spans about 31 hours. With two-level temporal boundaries, 6211 single actions are located, Fig. 4(a & b) present the duration distribution of action sequence videos and single action videos. The duration of action sequence videos is widely distributed between 33 and 186 seconds, and most actions last from 10 to 30 seconds. The distribution of the number of different types of action is shown in Fig. 4(c), where “Jumping jacks” takes up the highest proportion of takes (i.e., 14.2%), and the action takes up the fewest proportion of takes is “Sumo Squat” (i.e., 5.3%).
Action Quality Score Distribution. We also analyze the distribution of action quality scores for each type of action in Fig. 4(d). Here, the score for each single action video is calculated by averaging the quality scores annotated by all annotators.
Figure 4: Statistics of EgoExo-Fitness.
@@
Datasets Public. Scenarios Ego Exo Step Text Keypoint Comments Score Durations
@@
Ours Fitness 31h
@@ 3.5 Comparison with Related Datasets
Table 3: Comparison of the proposed EgoExo-Fitness and the concurrent EgoExo4D [23] dataset. Compared with EgoExo4D, the proposed EgoExo-Fitness collects videos of a new scenario and augments data with novel annotations. For fair comparison, scenarios of full-body actions are considered.
Datasets Scenarios Step Text Keypoint Comments Score Durations
@@
EgoExo-Fitness(Ours) Fitness 31h
Note that a recent proposed large-scale dataset Ego-Exo4D [23] also contains full-body (physical) action videos collected by synchronized ego-exo cameras. Besides, they both attend to how well an action is performed and propose novel corresponding annotations. What makes our dataset different from Ego-Exo4D lines in the following ways (also shown in Tab. 3):
  • New scenario is focused on. We focus on the scenario of natural fitness practicing and collect dynamic action sequence videos (containing 3 to 6 different actions). However, in Ego-Exo4D, a video is only associated with one type of action/task. This makes our dataset better suited for ego-exo full-body action studies on action boundaries and orders.
  • Novel annotations are provided. We provide text guidance and technical keypoints verification (both NOT included in Ego-Exo4D), which offer more intuitive and detailed identifications of what is done well and what can be improved in executions than expert commentary in Ego-Exo4D. Such annotations enable the pioneering exploration of interpretable action assessment.
  • Other unique characteristics. We also provide videos captured from two downward cameras for capturing more body details in movements, and annotations of two-level temporal boundaries to enable benchmark constructions.
For more details about the dataset comparisons, please refer to Appendix A4.

### v2

@@ Abstract
We present EgoExo-Fitness, a new full-body action understanding dataset, featuring fitness sequence videos recorded from synchronized egocentric and fixed exocentric (third-person) cameras. Compared with existing full-body action understanding datasets, EgoExo-Fitness not only contains videos from first-person perspectives, but also provides rich annotations. Specifically, two-level temporal boundaries are provided to localize single action videos along with sub-steps of each action. More importantly, EgoExo-Fitness introduces innovative annotations for interpretable action judgement–including technical keypoint verification, natural language comments on action execution, and action quality scores. Combining all of these, EgoExo-Fitness provides new resources to study egocentric and exocentric full-body action understanding across dimensions of “what”, “when”, and “how well”. To facilitate research on egocentric and exocentric full-body action understanding, we construct benchmarks on a suite of tasks (i.e., action classification, action localization, cross-view sequence verification, cross-view skill determination, and a newly proposed task of guidance-based execution verification), together with detailed analysis. Data and code are available at https://github.com/iSEE-Laboratory/EgoExo-Fitness/tree/main.
@@
Egocentric video dataset Full-body action understanding Fitness practising Interpretable action judgement
footnotetext: $\ddagger$: Project lead. $\dagger$: Equal key contributions. $\ast$: Corresponding authors.
Emails: {liym266, wanganlan}@mail2.sysu.edu.cn; mengjke@gmail.com; wszheng@ieee.org.
@@ 1 Introduction
By looking at the field of egocentric video understanding, we find that egocentric full-body action understanding is yet to be well explored due to the lack of datasets. Existing egocentric video datasets primarily focus on interactive actions like desktop works [37, 10, 11, 12, 59, 22, 2, 70, 28] (e.g., cooking and assembling) and daily interaction [23, 13, 54, 48, 62, 71] (i.e., interacting with daily objects or humans). The other branch of egocentric datasets [34, 80, 74, 25] mainly focuses on body pose estimation and reconstruction rather than understanding full-body action from other dimensions (e.g., verifying the consistency of action sequences and assessing the execution of action).
@@
  • Secondly, it provides two-level temporal boundaries to localize a single fitness action as well as sub-steps of each action.
  • More importantly, EgoExo-Fitness introduces annotations on interpretable action judgement, including technical keypoint verification, natural language comment, and quality score for each single action execution. To our knowledge, no previous dataset contains such annotations on action judgement.
Combining all of these, EgoExo-Fitness spans 32 hours with 1276 cross-view action sequence videos featuring more than 6000 single fitness actions. With synchronized ego-exo videos and rich annotations, EgoExo-Fitness provides new resources to study egocentric and exocentric full-body action understanding across dimensions of “what”, “when”, and “how well”.
@@
In summary, the contributions of our work are as follows: 1) We present EgoExo-Fitness, a new full-body action understanding dataset featuring fitness sequence videos recorded from synchronized ego- and exo-centric cameras; 2) We introduce rich annotations on EgoExo-Fitness, including two-level temporal boundaries and novel annotations of interpretable action judgement; 3) We construct benchmarks on five relevant vision tasks, including the newly introduced Guidance-based Execution Verification and extensive experimental analysis.
@@ 2.1 Revisiting Current Datasets
Full-Body Action Understanding datasets. Human body movements contain complex motion patterns and technical skills, presenting a series of challenges for Full-Body Action Understanding (FBAU). To address these challenges, datasets like NTU-RGB+D[60, 46], Human3.6M [30], Diving48 [41] and FineGym [61] are proposed to enable research on recognizing coarse-and-fine human full-body actions. Beyond recognition, datasets like Diving48-SV [55] and RepCount [27] are present to address tasks (e.g., Sequence Verification and Repetitive Action Counting) that require stronger temporal modeling ability. Note that technical full-body action videos (e.g., diving and vaulting) will reflect human skills. Hence, in recent years, datasets for Action Assessment, like AQA-7[51], FineDiving [73], LOGO[79], are introduced to study the subtle skill differences between action videos. Another branch of datasets [64, 19, 6] focuses on estimating or reconstructing 3D human poses from full-body action videos, achieving the development of Virtual Reality. Though great progress has been achieved, today’s full-body action understanding datasets mainly assume that human full-body action videos are captured by exocentric cameras. Such an assumption limits further exploration in more flexible settings. Moreover, some datasets (e.g., WEAR [5] and 1st-basketball [3]) propose to understand sports and fitness activities from egocentric viewpoints. However, these datasets are limited by their scales and task-specific annotations.
@@
Egocentric Video Understanding Datasets. Egocentric Video Understanding has great application values for AR/VR and Robotics. Most existing datasets focus on interactive actions: 1) tabletop activities in kitchen [10, 11, 12, 40, 66] or on a static working platform [37, 59, 22, 2, 57]; 2) daily activities interacting with daily objects [71, 38, 62, 48, 32, 28] or individuals [49, 58, 15]. Although recently proposed Ego4D [23] expands beyond interactive activities to a wider variety of daily activities, works on this branch of datasets rarely focus on egocentric full-body action understanding. Another branch of work aims to estimate full-body pose from egocentric videos, and several datasets [39, 80, 1, 69, 34] are released.
@@
Different from existing datasets, EgoExo-Fitness features synchronized egocentric and exocentric videos of full-body fitness actions and provides rich annotations (especially novel annotations of interpretable action judgement) for future research on understanding ego- and exo-centric full-body actions across the dimensions of “what”, “when”, and “how well”.
@@
It is worth noting that a concurrent large dataset, Ego-Exo4D [24], also contains ego-exo full-body (physical) action videos and annotations on how well an action is performed. EgoExo-Fitness still has its values: (1) It focuses on a novel scenario (i.e., natural fitness practising); (2) We provides novel annotations (e.g., technical keypoint verification), enabling studies on new tasks (e.g., Guidance-based Execution Verification). We will provide detailed comparisons between Ego-Exo4D and our work in Sec. 3.5 and Appendix A4.2.
@@ 2.2 Revisiting Relevant Tasks
Action Classification & Localization. As the fundamental tasks in video action understanding, action classification [44, 4, 9, 36] and temporal action localization [47, 78, 68, 18] are widely explored in previous work. In our work, we benchmark EgoExo-Fitness on action classification and localization to present the gap and characteristics across views in full-body action understanding.
Sequence Verification. Sequence Verification (SV) [55, 14, 26] is proposed to study the action order of sequential videos under a scenario that precise temporal annotations are not provided. Today’s benchmarks on SV rather focus on exocentric-SV (i.e., COIN-SV and Diving48-SV) or egocentric-SV (i.e., CSV). In this work, we present the first benchmark on cross-view sequence verification and provides extensive experimental analysis.
Action Assessment. Existing datasets on Action Assessment (or Skill Determination) are mainly based on videos from either ego-cameras [16, 21] or exo-cameras [73, 52, 51, 79, 72, 17, 76, 20, 77], which leads to single-view assessment ability. Also, today’s popular datasets only provide the annotations on action scores or paired rankings, which is unable existing work to directly explore the interpretability of the predicted results. To address the first issue, we introduce the first benchmark on Cross-View Skill Determination. For the second issue, we propose a novel task, Guidance-based Execution Verification, which aims to verify whether the execution of an action satisfies the given technical keypoints.
@@ 3 EgoExo-Fitness Dataset
In this section, we introduce the EgoExo-Fitness dataset in detail. We will start by describing the recording system in Sec. 3.1, data collection in Sec. 3.2, and annotations in Sec. 3.3. After that, we present the statistics and comparisons with related datasets in Sec. 3.4 and Sec. 3.5, respectively.
@@ 3.1 Recording System
We build a recording system for EgoExo-Fitness to capture action videos from egocentric and exocentric views. Fig. 2 shows the setups of our recording system. For egocentric video capturing, we design a headset with multiple cameras to record videos from forward and downwards views. Specifically, we use a GoPro to record the forward (i.e., Ego-M) view of participants and apply two Insta-Go3 cameras to record the left-downward (i.e. Ego-L) and right-downward (i.e., Ego-R) views. For the exocentric cameras, we locate them at the participants’ front (i.e., Exo-M), left-front (i.e., Exo-L), and right-front (i.e., Exo-R) sides and ensure they can record full-body actions completely. All cameras are synchronized manually using a timed event that is visible from them.
@@ 3.2 Action Sequence & Recording Protocols
Following FLAG3D [64] and HuMMan [6], we select 12 types of fitness actions based on various driving muscles (i.e., chest, abdomen, waist, hip, and whole body). All selected actions are listed in Tab. 1. In the remaining part of the paper, for convenience, we will use the abbreviation to present the actions. Furthermore, to enrich the temporal diversity of the recorded videos, we define 86 action sequences by randomly combining 3 to 6 different actions. For example, “starting with Push-ups, then Sit-ups, finally High Knee” is an action sequence with three fitness actions. For details of the action sequences, please refer to Appendix A2.
@@
Table 1: Recorded fitness actions. Abbr.: the abbreviation of the fitness action.
Action types Abbr. Action types Abbr. Action types Abbr.
1: Kneeling Push-ups KPU 5: Shoulder Bridge SB 10: Jumping Jacks JJ
2: Push-ups PU 6: Sit-ups SU 11: High Knee HK
3: Kneeling Torso Twist KTT 7: Leg Reverse Lunge LRL 12: Clap Jacks CJ
4: Knee Raise and Abdomi- KRAMC 8: Leg Lunge with Knee Lift LLKL
         nal Muscles Contract 9: Sumo Squat SS
@@ 3.3 Annotations
Interpretable Action Judgement. Our motivation for providing this series of annotations is two-fold. First, it is easy for human experts to compare an action video and the text guidance to conclude whether the actor followed the guidance or not and point out which technical keypoint in the text guidance is missed during the execution. However, such ability has rarely been studied in existing video-language answering and video-language retrieval works. Second, although Action Assessment [51, 72, 52, 73, 16, 17, 42] has been studied for many years, and great progress has been achieved, interpretable action assessment based on language annotations has never been explored due to the lack of well-collected dataset. Also, existing action assessment work is limited to ego-only or exo-only scenarios due to the collecting manner of the datasets.
To address these issues, we develop a web-based annotation tool for EgoExo-Fitness, and collect three categories of interpretable action judgement annotations (i.e., Technical keypoints Verification, Natural Language Comment, and Action Quality Score) step by step. We will introduce the details as follows.
@@
(1) Technical Keypoint Verification. A paragraph of text guidance on fitness action can be divided into several technical keypoints. By following the keypoints, one can achieve the goal of exercise while avoiding physical injury. In EgoExo-Fitness, we provide the verification annotations on the keypoints for located single actions in the following three steps. First, following FLAG3D [64], we provide a paragraph of text guidance for recorded actions. Second, we prompt LLM (e.g., ChatGPT) to separate the guidance into several keypoints. Third, we ask the annotators to verify an action by comparing the execution with the technical keypoints. Given an action and a technical keypoint, if the action satisfies the keypoint, an annotation of “True” will be noted (otherwise “False”).
@@
(3) Action Quality Score. Finally, the annotators are asked to score the actions from 1 to 5 (worst to best) based on the technical keypoint verifications and comments they have made.
@@ 3.4 Statistics
Figure 4: Statistics of the proposed EgoExo-Fitness dataset.
Number of recordings and Duration. EgoExo-Fitness collects 1276 cross-view action sequence videos from 86 action sequences, which spans about 32 hours. With two-level temporal boundaries, 6131 single actions are located, Fig. 4(a & b) present the duration distribution of action sequence videos and single action videos. The duration of action sequence videos is widely distributed between 33 and 186 seconds, and most actions last from 10 to 30 seconds. The distribution of the number of different types of action is shown in Fig. 4(c), where “Jumping jacks” takes up the highest proportion of takes (i.e., 12.6%), and the action takes up the fewest proportion of takes is “Sumo Squat” (i.e., 5.3%).
Action Quality Score Distribution. We also analyze the distribution of action quality scores for each type of action in Fig. 4(d). Here, the score for each single action is calculated by averaging the scores annotated by all annotators.
@@
Datasets Public. Scenarios Ego Exo Step Text Keypoint Comment Score Duration
@@
EgoExo-Fitness(Ours) Fitness 32h
@@ 3.5 Comparison with Related Datasets
Table 3: Comparison with the concurrent Ego-Exo4D [24] dataset. The proposed EgoExo-Fitness collects videos of a new scenario and augments data with novel annotations. For fair comparison, scenarios of full-body actions are considered.
Datasets Scenarios Step Text Keypoint Comment Score Duration
@@
EgoExo-Fitness(Ours) Fitness 32h
Note that a recent proposed large-scale dataset Ego-Exo4D [24] also contains full-body (physical) action videos collected by synchronized ego-exo cameras. Besides, they both attend to how well an action is performed and propose novel corresponding annotations. What makes our dataset different from Ego-Exo4D lines in the following ways (also shown in Tab. 3): (1) New scenario is focused on. We focus on the scenario of natural fitness practising and collect dynamic action sequence videos (containing 3 to 6 different actions). However, in Ego-Exo4D, a video is only associated with one type of action/task. This makes our dataset better suited for ego-exo full-body action studies on action boundaries and orders. (2) Novel annotations are provided. We provide text guidance and technical keypoints verification (both NOT included in Ego-Exo4D), which offer another branch of intuitive and detailed identifications of what is done well and what can be improved in executions than expert commentary in Ego-Exo4D. Such annotations enable the pioneering exploration of interpretable action assessment. (3) Other unique characteristics. We also provide videos captured from two downward ego-cameras for capturing more body details in movements, and annotations of two-level temporal boundaries to enable benchmark constructions.
For more details about dataset comparisons, please refer to Appendix A4.